### v1

@@ A Case Study on Scaffolding Exploratory Data Analysis for AI Pair Programmers Zhou and Li (2023) 2023
Designing PairBuddy—A Conversational Agent for Pair Programming Robe and Kuttal (2022) 2022
Developers’ Perspective on Today’s and Tomorrow’s Programming Tool Assistance: A Survey Kuang et al. (2024) 2024
@@
Discovering the Syntax and Strategies of Natural Language Programming with Generative Language Models Jiang et al. (2022) 2022
@@
Documentation Matters: Human-Centered AI System to Assist Data Science Code Documentation in Computational Notebooks Wang et al. (2022) 2022
Enhancing Code Completion with Implicit Feedback Jin et al. (2023) 2023
Evaluating Human-AI Partnership for LLM-based Code Migration Omidvar Tehrani and Anubhai (2024) 2024
@@
Expectation vs. Experience: Evaluating the Usability of Code Generation Tools Powered by Large Language Models Vaithilingam et al. (2022) 2022
Exploring Interaction Patterns for Debugging: Enhancing Conversational Capabilities of AI-assistants Chopra et al. (2024) 2024
Exploring the Learnability of Program Synthesizers by Novice Programmers Jayagopal et al. (2022) 2022
Exploring the Problems, Their Causes and Solutions of AI Pair Programming: A Study on GitHub and Stack Overflow Zhou et al. (2025) 2024
From “Ban It Till We Understand It” to “Resistance is Futile”: How University Programming Instructors Plan to Adapt as More Students Use AI Code Generation and Explanation Tools such as ChatGPT and GitHub Copilot Lau and Guo (2023) 2023
Generation Probabilities Are Not Enough: Uncertainty Highlighting in AI Code Completions Vasconcelos et al. (2024) 2024
@@
Github copilot in the classroom: learning to code with AI assistance Puryear and Sprint (2022) 2022
Grounded Copilot: How Programmers Interact with Code-Generating Models Barke et al. (2023) 2023
How Do Data Analysts Respond to AI Assistance? A Wizard-of-Oz Study Gu et al. (2023) 2024
How Novices Use LLM-Based Code Generators to Solve CS1 Coding Tasks in a Self-Paced Learning Environment Kazemitabaar et al. (2023) 2023
@@
How Readable is Model-generated Code? Examining Readability and Visual Inspection of GitHub Copilot Al Madi (2022) 2022
IDA: Breaking Barriers in No-code UI Automation Through Large Language Models and Human-Centric Design Shlomov et al. (2024) 2024
Identifying the Factors That Influence Trust in AI Code Completion Brown et al. (2024) 2024
Improving Steering and Verification in AI-Assisted Data Analysis with Interactive Task Decomposition Kazemitabaar et al. (2024) 2024
In-IDE Generation-based Information Support with a Large Language Model Nam et al. (2023a) 2024
Investigating and Designing for Trust in AI-powered Code Generation Tools Wang et al. (2024) 2024
@@
Investigating Explainability of Generative AI for Code through Scenario-based Design Sun et al. (2022) 2022
Investigating Interaction Modes and User Agency in Human-LLM Collaboration for Domain-Specific Data Analysis Guo et al. (2024) 2024
@@
Is GitHub Copilot a Substitute for Human Pair-Programming? An Empirical Study Imai (2022) 2022
Is GitHub’s Copilot as Bad as Humans at Introducing Vulnerabilities in Code? Asare et al. (2023) 2023
Ivie: Lightweight Anchored Explanations of Just-Generated Code Yan et al. (2024) 2024
Lost at C: A User Study on the Security Implications of Large Language Model Code Assistants Sandoval et al. (2023) 2023
Methodology for Code Synthesis Evaluation of LLMs Presented by a Case Study of ChatGPT and Copilot Ságodi et al. (2024) 2024
Multi-line AI-assisted Code Authoring Dunay et al. (2024) 2024
Non-Expert Programmers in the Generative AI Future Feldman and Anderson (2024) 2024
On the Design of AI-powered Code Assistants for Notebooks McNutt et al. (2023) 2023
On the Robustness of Code Generation Techniques: An Empirical Study on GitHub Copilot Mastropaolo et al. (2023) 2023
Performance, Workload, Emotion, and Self-Efficacy of Novice Programmers Using AI Code Generation Gardella et al. (2024) 2024
Practices and Challenges of Using GitHub Copilot: An Empirical Study Zhang et al. (2023b) 2023
Practitioners’ expectations on automated code comment generation Hu et al. (2022) 2022
Practitioners’ Expectations on Code Completion Wang et al. (2023a) 2023
Productivity Assessment of Neural Code Completion Ziegler et al. (2022) 2022
Prompt Sapper: A LLM-Empowered Production Tool for Building AI Chains Cheng et al. (2024b) 2024
Reading Between the Lines: Modeling User Behavior and Costs in AI-Assisted Programming Mozannar et al. (2024a) 2024
Significant Productivity Gains through Programming with Large Language Models Weber et al. (2024) 2024
Slide4N: Creating Presentation Slides from Computational Notebooks with Human-AI Collaboration Wang et al. (2023b) 2023
Spellburst: A Node-based Interface for Exploratory Creative Coding with Natural Language Prompts Angert et al. (2023) 2023
@@
Taking Flight with Copilot: Early insights and opportunities of AI-powered pair-programming tools Bird et al. (2022) 2023
@@
The Impact of AI on Developer Productivity: Evidence from GitHub Copilot Peng et al. (2023) 2023
@@
The Programmer’s Assistant: Conversational Interaction with a Large Language Model for Software Development Ross et al. (2023a) 2023
The RealHumanEval: Evaluating Large Language Models’ Abilities to Support Programmers Mozannar et al. (2024c) 2024
The Widening Gap: The Benefits and Harms of Generative AI for Novice Programmers Prather et al. (2024) 2024
Toward Effective AI Support for Developers: A survey of desires and concerns Khemka and Houck (2024) 2024
Towards Feature Engineering with Human and AI’s Knowledge: Understanding Data Science Practitioners’ Perceptions in Human&AI-Assisted Feature Engineering Design Zhu et al. (2024) 2024
@@
Towards More Effective AI-Assisted Programming: A Systematic Design Exploration to Improve Visual Studio IntelliCode’s User Experience Vaithilingam et al. (2023) 2023
Transforming Software Development: Evaluating the Efficiency and Challenges of GitHub Copilot in Real-World Projects Pandey et al. (2024) 2024
Trust in Generative AI among students: An Exploratory Study Amoozadeh et al. (2024) 2024
Using AI Assistants in Software Development: A Qualitative Study on Security Practices and Concerns Klemmer et al. (2024) 2024
@@
Using GitHub Copilot to Solve Simple Programming Problems Wermelinger (2023) 2023
Validating AI-Generated Code with Live Programming Ferdowsi et al. (2024) 2024
@@
When to Show a Suggestion? Integrating Human Feedback in AI-Assisted Programming Mozannar et al. (2024b) 2024
@@ 3 RESULTS
This section presents key findings from our review of the studies on human-AI interaction in software development. We analyze methodological trends, study contexts, AI tool design, AI’s impact on productivity and code quality, and emerging research directions. As described in Figure 1 and Section 2, we screened 254 papers, 100 of which were assessed for eligibility, with 89 included in the final review stage.

### v2

@@ 1 An Empirical Evaluation of GitHub Copilot’s Code Suggestions Nguyen and Nadi (2022)
5 Designing PairBuddy—A Conversational Agent for Pair Programming Robe and Kuttal (2022)
@@
6 Discovering the Syntax and Strategies of Natural Language Programming with Generative Language Models Jiang et al. (2022)
@@
7 Documentation Matters: Human-Centered AI System to Assist Data Science Code Documentation in Computational Notebooks Wang et al. (2022)
@@
8 Expectation vs. Experience: Evaluating the Usability of Code Generation Tools Powered by Large Language Models Vaithilingam et al. (2022)
9 Exploring the Learnability of Program Synthesizers by Novice Programmers Jayagopal et al. (2022)
@@
10 Github copilot in the classroom: learning to code with AI assistance Puryear and Sprint (2022)
@@
11 How Readable is Model-generated Code? Examining Readability and Visual Inspection of GitHub Copilot Al Madi (2022)
@@
12 Investigating Explainability of Generative AI for Code through Scenario-based Design Sun et al. (2022)
@@
13 Is GitHub Copilot a Substitute for Human Pair-Programming? An Empirical Study Imai (2022)
14 Practitioners’ expectations on automated code comment generation Hu et al. (2022)
15 Productivity Assessment of Neural Code Completion Ziegler et al. (2022)
@@ 13 Is GitHub Copilot a Substitute for Human Pair-Programming? An Empirical Study Imai (2022)
16 Taking Flight with Copilot: Early insights and opportunities of AI-powered pair-programming tools Bird et al. (2022)
17 “It’s Weird That it Knows What I Want”: Usability and Interactions with Copilot for Novice Programmers Prather et al. (2023)
18 A Case Study in Engineering a Conversational Programming Assistant’s Persona Ross et al. (2023b)
19 A Case Study on Scaffolding Exploratory Data Analysis for AI Pair Programmers Zhou and Li (2023)
20 A Mixed Reality Approach for Innovative Pair Programming Education with a Conversational AI Virtual Avatar Manfredi et al. (2023)
21 AI-Powered Chatbots and the Transformation of Work: Findings from a Case Study in Software Development and Software Engineering Süße et al. (2023)
22 Anticipating User Needs: Insights from Design Fiction on Conversational Agents for Computational Thinking Penney et al. (2023)
23 Case Study: Using AI-Assisted Code Generation In Mobile Teams Vasiliniuc and Groza (2023)
24 CoLadder: Supporting Programmers with Hierarchical Code Generation in Multi-Level Abstraction Yen et al. (2023)
25 Conversing with Copilot: Exploring Prompt Engineering for Solving CS1 Problems Using Natural Language Denny et al. (2023)
26 Copilot for Xcode: Exploring AI-Assisted Programming by Prompting Cloud-based Large Language Models Tan et al. (2023)
27 From “Ban It Till We Understand It” to “Resistance is Futile”: How University Programming Instructors Plan to Adapt as More Students Use AI Code Generation and Explanation Tools such as ChatGPT and GitHub Copilot Lau and Guo (2023)
28 Grounded Copilot: How Programmers Interact with Code-Generating Models Barke et al. (2023)
29 How Novices Use LLM-Based Code Generators to Solve CS1 Coding Tasks in a Self-Paced Learning Environment Kazemitabaar et al. (2023b)
30 In-IDE Generation-based Information Support with a Large Language Model Nam et al. (2023)
31 Is GitHub’s Copilot as Bad as Humans at Introducing Vulnerabilities in Code? Asare et al. (2023)
32 Lost at C: A User Study on the Security Implications of Large Language Model Code Assistants Sandoval et al. (2023)
33 On the Design of AI-powered Code Assistants for Notebooks McNutt et al. (2023)
34 On the Robustness of Code Generation Techniques: An Empirical Study on GitHub Copilot Mastropaolo et al. (2023)
35 Practices and Challenges of Using GitHub Copilot: An Empirical Study Zhang et al. (2023)
36 Practitioners’ Expectations on Code Completion Wang et al. (2023a)
37 Slide4N: Creating Presentation Slides from Computational Notebooks with Human-AI Collaboration Wang et al. (2023b)
38 Spellburst: A Node-based Interface for Exploratory Creative Coding with Natural Language Prompts Angert et al. (2023)
39 Studying the effect of AI Code Generators on Supporting Novice Learners in Introductory Programming Kazemitabaar et al. (2023a)
@@ 23 Case Study: Using AI-Assisted Code Generation In Mobile Teams Vasiliniuc and Groza (2023)
40 The Impact of AI on Developer Productivity: Evidence from GitHub Copilot Peng et al. (2023)
@@
41 The Programmer’s Assistant: Conversational Interaction with a Large Language Model for Software Development Ross et al. (2023a)
@@
42 Towards More Effective AI-Assisted Programming: A Systematic Design Exploration to Improve Visual Studio IntelliCode’s User Experience Vaithilingam et al. (2023)
@@
43 Using GitHub Copilot to Solve Simple Programming Problems Wermelinger (2023)
44 “It would work for me too”: How Online Communities Shape Software Developers’ Trust in AI-Powered Code Generation Tools Cheng et al. (2024a)
45 A Large-Scale Survey on the Usability of AI Programming Assistants: Successes and Challenges Liang et al. (2024)
46 A Study on Developer Behaviors for Validating and Repairing LLM-Generated Code Using Eye Tracking and IDE Actions Tang et al. (2024)
47 A Transformer-Based Approach for Smart Invocation of Automatic Code Completion de Moor et al. (2024)
48 A User-centered Security Evaluation of Copilot Asare et al. (2024)
49 AI Tool Use and Adoption in Software Development by Individuals and Organizations: A Grounded Theory Study Li et al. (2024)
50 An Analysis of the Costs and Benefits of Autocomplete in IDEs Jiang and Coblenz (2024)
51 An Empirical Study of Code Search in Intelligent Coding Assistant: Perceptions, Expectations, and Directions Liu et al. (2024)
52 An Empirical Study on Usage and Perceptions of LLMs in a Software Engineering Project Rasnayaka et al. (2024)
53 An Exploratory Study on Upper-Level Computing Students’ Use of Large Language Models as Tools in a Semester-Long Project Tanay et al. (2024)
54 Analyzing Prompt Influence on Automated Method Generation: An Empirical Study with Copilot Fagadau et al. (2024)
55 Ansible Lightspeed: A Code Generation Service for IT Automation Sahoo et al. (2024)
56 Are Prompt Engineering and TODO Comments Friends or Foes? An Evaluation on GitHub Copilot OBrien et al. (2024)
57 Can Developers Prompt? A Controlled Experiment for Code Documentation Generation Kruse et al. (2024)
58 Copilot Evaluation Harness: Evaluating LLM-Guided Software Programming Agarwal et al. (2024)
59 CoPrompt: Supporting Prompt Sharing and Referring in Collaborative Natural Language Programming Feng et al. (2024)
60 Defendroid: Real-time Android code vulnerability detection via blockchain federated neural network with XAI Senanayake et al. (2024)
61 Design Principles for Collaborative Generative AI Systems in Software Development Chen and Zacharias (2024)
62 Developers’ Perspective on Today’s and Tomorrow’s Programming Tool Assistance: A Survey Kuang et al. (2024)
63 Evaluating Human-AI Partnership for LLM-based Code Migration Omidvar Tehrani et al. (2024)
64 Exploring Interaction Patterns for Debugging: Enhancing Conversational Capabilities of AI-assistants Chopra et al. (2024)
65 How Do Data Analysts Respond to AI Assistance? A Wizard-of-Oz Study Gu et al. (2024)
66 IDA: Breaking Barriers in No-code UI Automation Through Large Language Models and Human-Centric Design Shlomov et al. (2024)
67 Identifying the Factors That Influence Trust in AI Code Completion Brown et al. (2024)
68 Improving Steering and Verification in AI-Assisted Data Analysis with Interactive Task Decomposition Kazemitabaar et al. (2024a)
69 Investigating and Designing for Trust in AI-powered Code Generation Tools Wang et al. (2024)
70 Investigating Interaction Modes and User Agency in Human-LLM Collaboration for Domain-Specific Data Analysis Guo et al. (2024)
71 Ivie: Lightweight Anchored Explanations of Just-Generated Code Yan et al. (2024)
72 Methodology for Code Synthesis Evaluation of LLMs Presented by a Case Study of ChatGPT and Copilot Ságodi et al. (2024)
73 Multi-line AI-assisted Code Authoring Dunay et al. (2024)
74 Non-Expert Programmers in the Generative AI Future Feldman and Anderson (2024)
75 Performance, Workload, Emotion, and Self-Efficacy of Novice Programmers Using AI Code Generation Gardella et al. (2024)
76 Prompt Sapper: A LLM-Empowered Production Tool for Building AI Chains Cheng et al. (2024b)
77 Reading Between the Lines: Modeling User Behavior and Costs in AI-Assisted Programming Mozannar et al. (2024a)
78 Significant Productivity Gains through Programming with Large Language Models Weber et al. (2024)
79 The RealHumanEval: Evaluating Large Language Models’ Abilities to Support Programmers Mozannar et al. (2024c)
80 The Widening Gap: The Benefits and Harms of Generative AI for Novice Programmers Prather et al. (2024)
81 Toward Effective AI Support for Developers: A survey of desires and concerns Khemka and Houck (2024)
82 Towards Feature Engineering with Human and AI’s Knowledge: Understanding Data Science Practitioners’ Perceptions in Human&AI-Assisted Feature Engineering Design Zhu et al. (2024)
83 Transforming Software Development: Evaluating the Efficiency and Challenges of GitHub Copilot in Real-World Projects Pandey et al. (2024)
84 Trust in Generative AI among students: An Exploratory Study Amoozadeh et al. (2024)
85 Using AI Assistants in Software Development: A Qualitative Study on Security Practices and Concerns Klemmer et al. (2024)
86 Validating AI-Generated Code with Live Programming Ferdowsi et al. (2024)
@@ 81 Toward Effective AI Support for Developers: A survey of desires and concerns Khemka and Houck (2024)
87 When to Show a Suggestion? Integrating Human Feedback in AI-Assisted Programming Mozannar et al. (2024b)
88 Exploring the Design Space of Cognitive Engagement Techniques with AI-Generated Code for Enhanced Learning Kazemitabaar et al. (2025)
89 Exploring the problems, their causes and solutions of AI pair programming: A study on GitHub and Stack Overflow Zhou et al. (2025b)
90 Generation Probabilities Are Not Enough: Uncertainty Highlighting in AI Code Completions Vasconcelos et al. (2025)
@@ 3 RESULTS
This section presents key findings from our review of the studies on in-IDE HAX. We analyze the contexts of the studies, the reported impact of AI on developers, the design of AI tools for coding, and the quality of the AI tools. The expanded corpus confirms the adequacy of the initial three-part framing (Sergeyuk et al., 2024) and adds finer subscopes within each dimension, but it does not introduce any new top-level dimensions. We also examine future work directions proposed in the reviewed literature and methodological patterns of the studies. As described in Figure 1 and Section 2, we screened a total of 256 papers, 100 of which were assessed for eligibility, with 90 included in the final review stage.
3.1 Context of studies
Figure 2: Paper IDs by Context.
The studies analyzed in this review represent two primary contexts for working with AI in software development: professional (71 out of 90) and educational (19/90) (see Fig. 2 for paper IDs). These contexts provide contrasting perspectives on the in-IDE integration of AI tools, with professional studies focusing on real-world applications and educational studies exploring their role in learning environments. Note that studies in educational settings may be underrepresented because our scope targets in-IDE contexts and students do not always work inside full IDEs during coursework. For a literature review of broader AI tools for coding in a strictly educational context, we refer the reader to  (Agbo et al., 2025; Ariza et al., 2025)
The majority of professional studies emphasize the practical utility of AI tools. Studies in this context address coding (17/71) and maintenance (8/71), with limited exploration of requirements gathering (1/71). However, 46 out of 71 studies do not specify a target SDLC stage, instead evaluating AI tooling for general programming use without stage-level framing, indicating a gap in contextual precision.
Educational studies primarily focus on pedagogical insights. These studies often investigate how students use AI tools to learn programming or solve computational problems. While a few studies address coding (6/16) and requirements formulation (1/16), most (13/16) do not specify the SDLC stage, aligning with the broader exploratory nature of educational research.
Figure 3: Paper IDs by Context and SDLC stage.
The distribution depicted on Figure 3, reveals notable gaps in the literature:
  • Lack of perspectives from people who do not use AI tools for development for a variety of reasons.
  • Limited number of educational studies highlighting the need for a deeper exploration of AI ’s impact on practical skills acquired during CS education, especially as developers constantly learn and re-learn programming practices in development environments.
Figure 4: Paper IDs by Types of tasks and Year.
Analysis of tasks assigned to the participants of studies reveals that they change over the three years. Early studies rely on smaller algorithmic exercises similar to LeetCode [^3] LeetCode: a resource for technical interviews materials and coding competitions https://leetcode.com/ -type programming tasks. Later studies introduce tasks that require data analysis, or a project-level use of AI. Data analysis tasks become more common over time. Studies add wrangling, plotting, and notebook workflows, which reflects the maturity of AI support in these environments and the relevance of studying such use. This broadening is also visible in the growth of the Unspecified category in Fig. 4, where authors evaluate AI assistance as general programming support without tying it to a single topic or SDLC stage. The variety of assigned tasks increases across years. The corpus now includes not only algorithmic coding and data analysis, but also documentation, testing, and a few cases of debugging and migration. This spread enables more precise claims about where AI helps, but also raises the need for clearer task descriptions to support comparison across studies.
A subset of studies conceptualizes programming with AI as a form of pair programming, where AI acts as a collaborative partner (Zhou et al., 2025b; Imai, 2022; Zhou and Li, 2023; Bird et al., 2022; Kazemitabaar et al., 2023a). This perspective highlights the evolving view of AI as not just a tool but a collaborator, representing a growing trend in co-creative programming workflows and suggesting new opportunities and challenges for both practitioners and educators.
Figure 5: Paper Counts and IDs by Scope.
3.2 Impact
Most prominent branch of research is dedicated to Impact of HAX in IDE with 74 out of 90 (see Fig. 5 for IDs) reviewed studies focusing on it. The increasing integration of AI-powered tools in IDEs has reshaped the way developers engage with coding, debugging, documentation, and software design (Peng et al., 2023; Weber et al., 2024). AI ’s role extends beyond simple automation, influencing productivity and software engineering practices, e.g.,  (Nguyen and Nadi, 2022; Vaithilingam et al., 2022). Although AI has undoubtedly enhanced productivity in some contexts, its overall effects remain highly context-dependent and nuanced. Studies consistently highlight the dual nature of AI’s impact, providing both benefits and challenges.
Multiple studies looked into the productivity aspect of coding with AI (13/74). Some of them have confirmed a notable boost in productivity when developers incorporate AI into their workflows. Potential benefits include a higher “ceiling” of productivity for experienced developers and faster onboarding for newcomers (Vasiliniuc and Groza, 2023). For instance, participants using GitHub Copilot completed implementation of an HTTP server in JavaScript up to 55.8% faster than control groups (Peng et al., 2023), while other investigators observed productivity gains ranging from 26% to 35% for complex tasks, across many files, and in proprietary contexts (Pandey et al., 2024). These improvements are partially attributed to fewer context switches and reduced boilerplate coding, enabling developers to offload repetitive tasks to AI while focusing on higher-level logic. However, gains are inconsistent: developers report that tasks involving proprietary or highly complex logic see only marginal benefits, as current assistance struggles with context-specific nuances of the real-world programming (Khemka and Houck, 2024).
Although AI-driven suggestions frequently accelerate development, they also require new verification efforts. When using AI tools, verifying suggestions, refining prompts, and reworking code generated by AI might take up to 50% of developers’ time (Mozannar et al., 2024a). This increased mental overhead arises because AI ’s output can be partially correct yet subtly flawed (Wermelinger, 2023), requiring careful review and, at times, stepwise interaction or re-prompting (Kazemitabaar et al., 2024a). Recent work proposes mechanisms that reduce this cost, showing that highlighting tokens by predicted edit likelihood improves speed and directs attention to problematic regions (Vasconcelos et al., 2025). Such extra validation often mitigates the risk of blindly accepting erroneous suggestions, i.e., phenomenon known as “automation bias” (Al Madi, 2022).
Studies focusing on novice programmers reveal both benefits and pronounced risks. Across classroom and lab settings, AI assistance tends to speed novice work and reduce perceived workload while short-term learning remains intact when use is structured (Kazemitabaar et al., 2023a). Beginners often solve programming assignment quicker with AI’s help (Puryear and Sprint, 2022; Rasnayaka et al., 2024; Tanay et al., 2024). Time-pressured studies report lower mental workload and steady self-efficacy with assistance, with additional time using AI relating to better task outcomes (Gardella et al., 2024). However, beginners can over-trust AI, losing sight of important foundational concepts (Prather et al., 2024). In certain cases, when system suggestions are misleading, novices drift away from correct solutions, as they lack the expertise to detect subtle errors (Zhou and Li, 2023). Structured prompts and explanatory interfaces can help novices understand AI outputs better, fostering safer learning pathways (Kruse et al., 2024). A hybrid strategy that mixes authoring and modification is associated with better results in self-paced CS1 tasks, and prompting is a learnable activity that raises solve rates when taught explicitly (Kazemitabaar et al., 2023b; Denny et al., 2023).
3.2.1 Attitude
A subset of studies (11/74) examines user attitude toward AI assistance in IDEs, with trust as the central construct. Trust is shaped by both the properties of the suggestions and the expertise of the user (Amoozadeh et al., 2024; Brown et al., 2024). High-quality and context-relevant suggestions raise trust, while inconsistency and unnecessary complexity lower it. Layers of outputs that help users judge suggestions, such as brief explanations, usage statistics, and control over scope, are proposed as practical supports for trust formation (Wang et al., 2024). Context also matters. Trust drops for high-stakes or production-related work and for complex or open-ended tasks, and it rises for routine work or proof-of-concept code (Wang et al., 2024). In addition, developers treat suggestions in test files more cautiously than those in production files, and they are less likely to accept code completion suggestions when writing tests (Brown et al., 2024).
Attitude relates to reliance. Over-reliance appears when users accept AI output without sufficient checking, and under-reliance appears when they dismiss correct output because justification feels costly. Novices often overestimate the reliability of AI output and invest less time in verification for algorithmic problems (Al Madi, 2022), so they tend to over-rely on AI. In educational settings, reliance and trust vary widely and depend on perceived performance of the system and task context, which warrants explicit trust calibration rather than blanket bans (Amoozadeh et al., 2024; Brown et al., 2024; Wang et al., 2024; Lau and Guo, 2023). Students also renegotiate authorship and autonomy when assistants anticipate intent, which calls for assessment that values reasoning traces and verification (Prather et al., 2023). Moreover, novices struggle to form accurate mental models of synthesizers, so interfaces should expose system state and allow incremental control (Jayagopal et al., 2022). Professional developers sometimes under-rely and reject correct outputs that lack a clear rationale (Sun et al., 2022). Tooling can mediate both patterns. Live programming environments that surface runtime values reduce the effort of review and support more calibrated acceptance decisions (Ferdowsi et al., 2024).
3.3 Design
To account for the changes introduced by AI integration in IDEs, a focused body of research has developed on the Design of HAX in IDE. Among the reviewed studies, 28 out of 90 (see Fig. 5 for IDs) investigate forms of AI integration into IDEs, with most of them (22/28) explicitly examining how these forms of integration impact users, and 17 of them exploring design principles.
The research presents two primary paradigms of AI-assisted interaction inside IDEs: autocompletion-based assistance and conversational agents. Autocompletion interfaces provide short, low-friction code completions, enabling developers to work seamlessly with inline suggestions. These approaches minimize cognitive load and support rapid decision-making without disrupting workflow (Vaithilingam et al., 2023; Yen et al., 2023). In contrast, conversational interactions (often embodied in chat-based assistants or virtual avatars) facilitate higher-level reasoning, iterative refinement, and complex problem-solving (Ross et al., 2023a; Robe and Kuttal, 2022; Gu et al., 2024). Rather than being opposing paradigms, these interaction modes are complementary (Weber et al., 2024), with studies advocating for hybrid models that seamlessly integrate autocompletion with interactive, context-sensitive conversations. In addition, some studies experiment with alternative modalities of interaction that go beyond these two dominant paradigms. For example, mixed-reality interfaces (Manfredi et al., 2023) introduce conversational AI in augmented pair-programming environments, leveraging immersive learning experiences.
Due to the conversational nature of interaction with AI in IDE, prompting became a prominent object of investigations with 7 out of 28 studies focusing on this aspect. Across reviewed studies, prompt engineering is positioned as a strategic skill that shapes how developers and AI systems collaborate, affecting both individual productivity and team-based workflows (Kruse et al., 2024; Yen et al., 2023). There are claims that prompting should be taught as a skill, with scaffolds for decomposition, short rationales before revealing solutions, and reusable prompt solution pairs delivered in step-by-step dialogues  (Kazemitabaar et al., 2025). A central challenge in AI-assisted development is ensuring that AI-generated code aligns with user intent (Fagadau et al., 2024; Denny et al., 2023; OBrien et al., 2024). Therefore, studies investigating prompt formulation demonstrate that structured and well-contextualized prompts lead to more accurate and reliable output, while excessive detail can degrade performance. As AI tools become embedded in professional and educational workflows, collaborative prompt engineering is also gaining importance (Feng et al., 2024; Cheng et al., 2024b). Consequently, authors introduce frameworks for shared prompt refinement, knowledge transfer, and workflow optimization, suggesting the need for modularity and re-usability in AI interactions.
For education, in-IDE HAX should shift learners from code consumption to deliberate practice. Interfaces need small, purposeful friction before code adoption, which would improve transfer and calibration without added load (Kazemitabaar et al., 2025). Contribution of AI should be staged and gated by quick checks (Robe and Kuttal, 2022; Kazemitabaar et al., 2025). Embodied variants can sustain engagement with the task when a studying partner is absent, as mixed reality pair-programming with a conversational avatar ties feedback to editor context (Manfredi et al., 2023). Finally, to reduce over-reliance, systems should make performance on similar tasks completed without assistance visible for self-assessment (Kazemitabaar et al., 2025).
Beyond interaction paradigms, 17 out of 28 studies explore design principles for AI-powered systems inside IDEs, for instance (Chen and Zacharias, 2024; Robe and Kuttal, 2022; Jayagopal et al., 2022; Cheng et al., 2024a; Yen et al., 2023; Shlomov et al., 2024; Guo et al., 2024; McNutt et al., 2023; Wang et al., 2023b; Feng et al., 2024; Gu et al., 2024; Angert et al., 2023; Kazemitabaar et al., 2024a). These principles guide how AI should be embedded into developer workflows and environments while maximizing usability, reliability, and efficiency:
  • Context Awareness: LLM-powered assistance significantly boosts task completion rates when it can effectively incorporate a developer’s code as context (Nam et al., 2023; Gu et al., 2024).
  • Explainability & Transparency: Tools should provide context-aware explanations and highlight uncertain AI-generated code (Vasconcelos et al., 2025; Yan et al., 2024; Sun et al., 2022; Wang et al., 2024; Vaithilingam et al., 2023) to improve user trust.
Among implementations of AI for IDEs, GitHub Copilot is by far the most studied, with 36 out of 90 works investigating it. This concentrated attention on a single tool suggests a risk of overgeneralization, where findings may not fully translate to other AI-powered coding assistants, limiting broader design and implementation insights.
3.4 Quality
Investigating and ensuring the Quality of AI-assisted code generation is a central concern of 19 out of 90 papers in the reviewed literature (see Fig. 5 for IDs). Across studies, quality is examined from multiple perspectives, including code correctness, security, and readability.
One recurring theme is the trade-off between efficiency and correctness. AI-assisted coding significantly accelerates development workflows, but at the cost of increased susceptibility to subtle errors, security vulnerabilities, and reduced maintainability (Sandoval et al., 2023; Tang et al., 2024; Pearce et al., 2022). Several studies suggest that while AI-generated code often appears syntactically valid, it can contain logical flaws, insecure patterns, or deviations from best practices (Asare et al., 2024; Pearce et al., 2022). As a result, effective quality control mechanisms, such as automated verification, static analysis, and model-aware debugging tools, are emphasized as critical to responsible adoption (Agarwal et al., 2024).
Beyond correctness, readability and maintainability emerge as key aspects of AI-generated code quality. Although AI can generate complex solutions rapidly, its output is not always optimized for human understanding. Studies highlight that AI-generated code may be less readable and comprehensible due to overly concise structures, unconventional variable naming, or lack of meaningful comments (Al Madi, 2022). This raises concerns about long-term maintainability, particularly in team-based software development workflows, where code legibility is as important as functional correctness. Research suggests that integrating AI-powered refactoring and explanation tools can help mitigate these challenges, enhancing both short-term efficiency and long-term code quality (Yan et al., 2024).
Studies additionally pose security as an additional concern: up to 36% of vulnerabilities in AI-assisted code originate from the LLMs, pointing out the risks of replicating insecure patterns from training data (Sandoval et al., 2023), emphasizing the need for cautious oversight of AI suggestions.
Overall, while LLM-powered tools can improve the software development process, maintaining consistently high-quality code remains a challenge. Ensuring reliability requires better alignment with developers’ needs and robust verification mechanisms (Tang et al., 2024; Agarwal et al., 2024). Studies highlight the need for strategies such as filtering suboptimal suggestions, improving contextual prompting through docstrings and function names (Mozannar et al., 2024c; de Moor et al., 2024). Moreover, research calls for adaptive personalisation of code suggestions to support the overall quality of AI-generated code. In practice, this means learning per-developer and per-task policies for when to suggest, which modality to use, how much code to propose, and how much context to include in the prompt (Mozannar et al., 2024c; de Moor et al., 2024).
Quality of code in AI-assisted development is not just a technical issue but also a human and organizational challenge. Research highlights that quality outcomes are influenced by how developers interact with AI tools: whether they verify AI-generated code, how they prompt AI models, and whether they develop strategies for AI-assisted debugging (Weber et al., 2024; Mozannar et al., 2024c; Sandoval et al., 2023; Al Madi, 2022; de Moor et al., 2024). Best practices for improving AI-assisted development quality extend beyond tool design and include education, training, and overall workflow adaptation.